Juan invests 30,000 at an annual rate of 2%. How many years (approximately the closest) will be spent to double his investment if the interest is compounded monthly, then continuously ?; Compare results.
Answer #1

compounded monthly In this case p= 30000, n=12, P=2%. = 2 = 0.02 and A= 30000 X 2 = 6 mo let t years will be spend to doubleCompounded continuously In This case A= pert 6.02.t = 60000 - 30000 e = 2 = e.o2t => 1og2 - (loge) (0.02) * to log 2 = 15 Y A